A rectangle is 72 sq.m and its perimeter is 34 m.find its length and breadth?
Answer by LinnW(1048)

set W = width
set L = length
L*W = 72
2W + 2L = 34
add -2L to each side
2W = 34 - 2L
divide each side by 2
W = 17 - L
substitute ( 17 - L ) for W in L*W = 72
L*( 17 - L ) = 72
17L - L^2 = 72
add -72 to each side
17L - L^2 - 72 = 0
multiply each side by -1
L^2 - 17L + 72 = 0
(L - 8)(L - 9) = 0
So L = 8 or L = 9
Suppose L = 9
Substitute 9 for L in 2W + 2L = 34
2W + 2(9) = 34
2W + 18 = 34
add -18 to each side
2W = 16
W = 8
So length is 9 and width is 8